# Adversarial Network Imagination: Causal LLMs and Digital Twins for Proactive Telecom Mitigation

### v1

@@ 1 Introduction
Digital twin technologies have emerged as a promising paradigm for modeling, monitoring, and optimizing complex networked systems, particularly in 5G and beyond networks (Ahmed et al., 2022; Ayoubi and others, 2019; Tao and Qi, 2019). Prior work has leveraged digital twins to simulate network topology, routing behavior, and resource utilization for offline diagnosis and planning (Ahmed et al., 2022; De Oliveira et al., 2022; Zhang et al., 2021). However, these approaches typically rely on manually crafted or historically observed failure scenarios. Similarly, fault-injection frameworks simulate predefined disruptions (Schölzel et al., 2018), but struggle to capture rare, coordinated, or cascading failures arising from complex dependency structures.
Causal modeling and knowledge graphs have been applied to represent dependencies among network components and improve fault localization (Pearl and Mackenzie, 2018; Kaur and Laliotis, 2021). While effective for diagnosis and explanation, these methods are largely non-generative and do not support the proactive exploration of counterfactual or adversarial failure scenarios. As a result, existing systems lack a principled mechanism to systematically imagine and evaluate novel failure patterns before they manifest in real networks.
@@ 2.1 Causal Modeling and Knowledge Graphs
Causal modeling provides a principled framework for representing dependencies and reasoning about interventions and counterfactuals (Pearl and Mackenzie, 2018). Recent work has explored causal representation learning and the integration of causal structure with machine learning models. Knowledge graphs offer a complementary abstraction to encode structured relationships and have been applied to fault diagnosis, dependency analysis, and root-cause localization in large-scale systems and networks (Kaur and Laliotis, 2021; Zhang et al., 2021).
In networks and systems domains, causal and graph-based models have been used to improve fault localization and explainability (Ayoubi and others, 2019; Ahmed et al., 2022). However, these approaches typically rely on predefined failure patterns or historical observations and lack generative mechanisms to actively explore novel or adversarial failure scenarios. Our work differs by leveraging explicit causal structure to guide generative scenario synthesis under structured dependency constraints.
@@ 2.2 LLMs for Reasoning and Structured Generation
Large language models have demonstrated strong reasoning and multi-step generation capabilities (Wei et al., 2022; Chowdhery et al., 2022), with techniques such as chain-of-thought prompting, tool use, and constrained decoding improving logical consistency and interpretability (Yao et al., 2023; Schick et al., 2023). Recent work has further explored incorporating external knowledge, constraints, or feedback signals to guide generation (Madaan et al., 2023; Bai et al., 2022).
@@ 3 Background and Problem Definition
Failure propagation in large-scale networked systems has been widely studied using graph-based and probabilistic models, with prior work examining cascading outages and dependency-driven disruptions (Meng et al., 2017). Related research in network tomography has explored inferring internal failures from end-to-end measurements, but such approaches often struggle to capture complex multi-hop dependencies (Nguyen and Thiran, 2016).
@@ 4.3 Knowledge Graph
The Knowledge Graph encodes the structural and functional dependencies of the telecom network, including topology links, routing paths, shared resources (e.g., CPU, memory, buffer capacities), and historical incidents. It provides a structured context for the Causal LLM to ensure that generated failures obey real-world constraints, such as component reachability, service dependencies, and load-sharing relationships. The graph also supports causal reasoning, allowing the LLM to infer indirect effects and cascading failure patterns that may emerge from primary disruptions. The causal modeling approach adopted in this work is inspired by foundational principles of causal inference, which emphasize reasoning about interventions and counterfactual outcomes rather than surface-level correlations (Pearl and Mackenzie, 2018).
@@ 6 Safety and Ethical Considerations
Digital twin deployments in telecom networks raise important security and resilience considerations (Salman et al., 2022). To mitigate operational risk, all failure scenarios are generated and executed exclusively within a controlled simulation environment; no actions are applied to live production networks. Knowledge graph representations and LLM inputs are fully anonymized to protect sensitive infrastructure and customer information (Gai et al., 2018). While the proposed framework enables systematic exploration of adversarial failure scenarios, all generated outputs are intended to support analysis and planning rather than automated decision-making, and should be interpreted under human oversight.

### v2

@@ 1 Introduction
Digital twin technologies have emerged as a promising paradigm for modeling, monitoring, and optimizing complex networked systems, particularly in 5G and beyond networks (Rodrigo et al., 2023; Kaur et al., 2021; Tao and Qi, 2019). Prior work has leveraged digital twins to simulate network topology, routing behavior, and resource utilization for offline diagnosis and planning (Rodrigo et al., 2023; Mozo et al., 2022; Chou and Jiang, 2021). However, these approaches typically rely on manually crafted or historically observed failure scenarios. Similarly, fault-injection frameworks simulate predefined disruptions, but struggle to capture rare, coordinated, or cascading failures arising from complex dependency structures.
Causal modeling and knowledge graphs have been applied to represent dependencies among network components and improve fault localization (Pearl, 2018; Peng et al., 2023). While effective for diagnosis and explanation, these methods are largely non-generative and do not support the proactive exploration of counterfactual or adversarial failure scenarios. As a result, existing systems lack a principled mechanism to systematically imagine and evaluate novel failure patterns before they manifest in real networks.
@@ 2.1 Causal Modeling and Knowledge Graphs
Causal modeling provides a principled framework for representing dependencies and reasoning about interventions and counterfactuals (Pearl, 2018). Recent work has explored causal representation learning and the integration of causal structure with machine learning models. Knowledge graphs offer a complementary abstraction to encode structured relationships and have been applied to fault diagnosis, dependency analysis, and root-cause localization in large-scale systems and networks (Peng et al., 2023; Chou and Jiang, 2021).
In networks and systems domains, causal and graph-based models have been used to improve fault localization and explainability (Kaur et al., 2021; Rodrigo et al., 2023). However, these approaches typically rely on predefined failure patterns or historical observations and lack generative mechanisms to actively explore novel or adversarial failure scenarios. Our work differs by leveraging explicit causal structure to guide generative scenario synthesis under structured dependency constraints.
@@ 2.2 LLMs for Reasoning and Structured Generation
Large language models have demonstrated strong reasoning and multi-step generation capabilities (Wei et al., 2022; Chowdhery et al., 2023), with techniques such as chain-of-thought prompting, tool use, and constrained decoding improving logical consistency and interpretability (Yao et al., 2022; Schick et al., 2023). Recent work has further explored incorporating external knowledge, constraints, or feedback signals to guide generation (Madaan et al., 2023; Bai et al., 2022).
@@ 3 Background and Problem Definition
Failure propagation in large-scale networked systems has been widely studied using graph-based and probabilistic models, with prior work examining cascading outages and dependency-driven disruptions. Related research in network tomography has explored inferring internal failures from end-to-end measurements, but such approaches often struggle to capture complex multi-hop dependencies (Lawrence et al., 2006).
@@ 4.3 Knowledge Graph
The Knowledge Graph encodes the structural and functional dependencies of the telecom network, including topology links, routing paths, shared resources (e.g., CPU, memory, buffer capacities), and historical incidents. It provides a structured context for the Causal LLM to ensure that generated failures obey real-world constraints, such as component reachability, service dependencies, and load-sharing relationships. The graph also supports causal reasoning, allowing the LLM to infer indirect effects and cascading failure patterns that may emerge from primary disruptions. The causal modeling approach adopted in this work is inspired by foundational principles of causal inference, which emphasize reasoning about interventions and counterfactual outcomes rather than surface-level correlations (Pearl, 2018).
@@ 6 Safety and Ethical Considerations
Digital twin deployments in telecom networks raise important security and resilience considerations (Wang et al., 2023). To mitigate operational risk, all failure scenarios are generated and executed exclusively within a controlled simulation environment; no actions are applied to live production networks. Knowledge graph representations and LLM inputs are fully anonymized to protect sensitive infrastructure and customer information (Gai et al., 2016). While the proposed framework enables systematic exploration of adversarial failure scenarios, all generated outputs are intended to support analysis and planning rather than automated decision-making, and should be interpreted under human oversight.